# Click Without Compromise: Online Advertising Measurement via Per User Differential Privacy

### v1

@@ 4.4. Ads-BPC Algorithm
Given the optimization problem in Problem (5), the optimal objective function value is $J=\left(\sum_{i=1}^{n}\sqrt{(n-i+1)r^{2}_{i}}\right)^{2}/(2\rho)$ with the optimal solution being $\sigma^{2}_{i}=\sqrt{(Jr^{2}_{i})/(2\rho(n-i+1))}$.
Note that the optimal noise distribution is non-i.i.d. distribution. Before introducing the noise addition algorithm, we outline our data truncation process, which limits each user’s contributions per day. For a fixed publisher, the truncation process operates at each timestamp $k$, tracking conversions ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{k}$ and their corresponding weights $w_{k}$. These weights indicate whether the conversion is attributed to impressions belonging to the given publisher. Specifically, $w_{k}=1$ if ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{k}$ is attributed to impressions from the given publisher, and $w_{k}=0$ if ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{c}}}_{k}$ is attributed to impressions from other publishers.
We maintain a dictionary, user_contribution, to aggregate the sum of attribution weights for each user. If a user’s sum of attribution weights exceeds a pre-specified threshold $r$, only the first $r$ sums of attribution weights will be retained in the dataset; any subsequent weights will be truncated. The detailed algorithm is presented in the Appendix 8.6.
Algorithm 1 outlines the process for releasing the user-level differentially private sum of attribution weights for a single publisher each day. In Line 1, we calculate the scale parameter using Theorem 4.5. Subsequently, in Line 1, we introduce standard Gaussian noise to the scaled version of the truncated ad measurements. This step is equivalent to adding noise from a $N(0,\sigma_{i}^{2})$ distribution to the truncated ad measurements. Finally, Line 1 reconstructs the total sum of attribution weights from day 1 to day $i$ and releases this private information.
Input: Number of days $n$, privacy budget $\rho$, the maximum user contributions at day $i$ is $r_{i}$, the truncated ads measurement on day $i$ is $x[i]$.
Output: The noisy ads measurement from day 1 to day $i$.
1 $J=\left(\sum_{i=1}^{n}\sqrt{(n-i+1)r^{2}_{i}}\right)^{2}/(2\rho)$
2 for $i=1\to n$ do
3       $\sigma^{2}_{i}=\sqrt{(Jr^{2}_{i})/(2\rho(n-i+1))}$
@@
4       $y_{i}=\frac{1}{\sigma_{i}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]+z_{i}$, $\quad z_{i}\sim N(0,1)$
5       Release $\sum_{t=1}^{i}\sigma_{t}*y_{t}$
6
Algorithm 1 Ads-BPC
4.5. Sanity Check for Proposed Method
Under the same setting of the small-scale numerical experiments in Table 7, we compare the optimal objective function values of the solver in (choquette2022multi, ) and our method. Table 8 shows that our method achieves optimality up to a difference of $10^{-2}$, which may come from numerical error.
4.6. Extension for Multiple Publishers
@@
Suppose there are $k$ publishers, the data vector for publisher $i$ is ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}_{i}\in\mathbb{R}^{n\times 1}$. Let $\mathbf{X}\in\mathbb{R}^{n\times k}$ be the concatenation of these $k$ data vectors ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}_{i}$. Similar to Section 4.3, the goal is to find a private answer to $\mathbf{W}\mathbf{X}$. We factorize the workload matrix $\mathbf{W}=\mathbf{L}\mathbf{B}$, where $\mathbf{B}=diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}})$ and $\mathbf{L}=\mathbf{W}\mathbf{B}^{-1}$. Let $\mathbf{Z}\sim MVG_{n,k}(0,\mathbf{I}_{n},\mathbf{I}_{k})$, where $MVG_{n,k}(0,\mathbf{I}_{n},\mathbf{I}_{k})$ is a Matrix Gaussian Distribution (chanyaswad2018mvg, ; yang2021improved, ), then the mechanism is
(6) ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}(\mathbf{X})=\mathbf{L}(\mathbf{B}\mathbf{X}+\mathbf{Z})=\mathbf{W}\mathbf{X}+\mathbf{L}\mathbf{Z}.$
@@
Theorem 4.6 (Sesitivity of Mechanism 6).
The sensitivity of the mechanism 6 under the definition of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ is
(7) $\Delta_{2}(f)=$ $\max_{(\mathbf{X},\tilde{\mathbf{X}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}}\|\mathbf{B}\mathbf{X}-\mathbf{B}\tilde{\mathbf{X}}\|_{F}=$ $\sqrt{\frac{r_{1}^{2}}{\sigma_{1}^{2}}+\cdots+\frac{r_{n}^{2}}{\sigma_{n}^{2}}}.$
Then similar to Theorem 4.4, we can show that the mechanism 6 also satisfies $\rho$-zCDP,
@@
Let $\mathbf{W}\in\mathbb{R}^{n\times n}$ be the prefix-sum matrix 4.1. Let $\sigma_{1},\sigma_{2},$ $\cdots\sigma_{n}$ be the solution to problem 5, $\mathbf{B}=diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}})$ and $\mathbf{L}=\mathbf{W}\mathbf{B}^{-1}$. Let $\mathbf{Z}\sim MVG_{n,k}(0,\mathbf{I}_{n},\mathbf{I}_{k})$, then ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}(\mathbf{X})$ = $\mathbf{L}(\mathbf{B}\mathbf{X}+\mathbf{Z})=\mathbf{W}\mathbf{X}+\mathbf{L}\mathbf{Z}$ satisfies $\rho$-zCDP.
Table 8. Comparison between the optimal objective function value of different methods.
$n$ Solver Our Method $n$ Solver Our Method
4 950.55 950.56 10 11365.32 11365.32
6 2400.11 2400.12 12 19441.07 19441.07
8 5193.02 5193.03 14 32607.31 32607.32
5. Discussion
In our specific context, it’s important to note that two data streams, denoted as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ and $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$, are considered neighbors when one can be derived from the other through the substitution of one user’s data. Notably, this solution can be derived from the traditional definition of neighboring datasets, where ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ and $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$ are regarded as neighbors when one can be obtained from the other by either adding or removing one user’s data. This transformation arises due to the possibility of utilizing twice the privacy budget: first for removing a user’s data and then for adding another user’s data.
In this paper, we primarily focus on rule-based attribution logics, including LTA, FTA, UNI, and other widely used multi-touch attribution methods. These methods assign attribution weights based on each user’s data. Methods that incorporate data across multiple users, such as incremental value heuristics and Shapley value (SV) methods, are left for future work.
It’s essential to recognize that our approach assumes the availability of upper bounds for user contributions, denoted as $r_{i}$. If a user surpasses the limit $r_{i}$ with records on day $i$, only the initial $r_{i}$ records will be retained in the dataset, while the excess will be truncated. Determining the appropriate values for $r_{i}$ entails navigating the bias-variance trade-off. Opting for a small $r_{i}$ leads to high bias, whereas a large $r_{i}$ introduces increased variance. Alternatively, a portion of the privacy budget could be allocated for examining the dataset and approximating the values of the upper bounds, $r_{i}$. However, we defer this task to future research, as the techniques involved in this estimation process are orthogonal to our current method.
Furthermore, it’s worth noting that our method offers robustness against malicious inputs exceeding the specified cap for individual contributions. This safeguard ensures that a single report cannot compromise the integrity of a query’s outcome by introducing unbounded trigger values.
6. Experiments
Throughout the experiments, we assume the length of the data vector is $n=len({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})$, that is, we need to report the accumulated attribution weights for $n$ days. For each dataset, we assume the global sensitivity is given as $GS$. In the case where the global sensitivity is unknown, we can set a reasonably large $GS$ as a threshold. Any user with more than $GS$ records in the dataset will only have the first $GS$ records retained, and any additional records will be removed.
On each day, we collect the relative error, then we report their median and 90% maximum (90th quantile). We will also report the root mean squared error (RMSE). For experiment with more than 1 publisher, we report the median error across all publishers. We repeat our experiments 10 times and report the average results. All experiments are conducted on a desktop with an 8-core 16-thread 3.6GHz Intel i9-9900K CPU and 32 GB memory.
6.1. Baseline Methods
We compare our algorithm (named Ads-BPC) with other baselines. Unless otherwise specified, we allocate a total privacy budget of $\rho_{total}=1$. We briefly introduce each baseline below and provide detailed descriptions in Appendix 8.7.

### v2

@@ 4.4 Estimate User Contribution Bound
Algorithm PrivateQuantile 2 satisfies $\rho_{2}$-zCDP with privacy paramter $\rho_{2}=\min(\frac{1}{8}\epsilon^{2},\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\epsilon)$.
4.5 Track the Change of Contribution Bound
If we run the private quantile algorithm every day, the privacy cost will be significant. We observe that the user contribution bound is stable most of the time, which means that the everyday estimate for the quantile is not necessary. If we keep the bound as a constant and only update it when there’s a significant change, it’s possible to save a large amount of privacy budget. In case when the contribution bound changes a lot in one day, we use Sparse Vector Technique (SVT) [34, 58, 18] to track such a change. We only need to report the user contribution bound when such a change is detected.
1:  Input: Query $q_{i}$; Privacy budget $\epsilon$; Number of reports count; Maximum number of reports $k$; Noisy threshold $\tilde{T}$; True threshold $T$.
12:  Output:
3:  update = False.
4:  if $\text{count}<k$ then
5:     {Stop if maximum reports reached.}
6:     Get noisy query: $\tilde{q}_{i}=q_{i}+\text{Lap}(4k/\epsilon)$.
7:     if $\tilde{q}_{i}>\tilde{T}$ then
8:        update = True.
9:        $\text{count}=\text{count}+1$.
10:        $\tilde{T}=T+\text{Lap}(2/\epsilon)$.
11:     end if
212:  end if
13:  Return (update, count, $\widetilde{T}$)
Algorithm 3 CheckUpdate
1:  Input: Dataset $D_{i}$.
2:  Parameter: Privacy budget $\epsilon$; Maximum number of reports $k$; Threshold $T^{\uparrow},T^{\downarrow}$; Scales ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\uparrow},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow}$
3:  Output: The user contribution bound $r_{i}$.
14:  Global Variable: Noisy threshold $\widetilde{T}^{\uparrow},\widetilde{T}^{\downarrow}$; The number of reports $\text{count}^{\uparrow},\text{count}^{\downarrow}$; Previous bounds: bound_list;
25:  Split $\epsilon$ into $\epsilon^{\uparrow}=\epsilon^{\downarrow}=\frac{\epsilon}{2}$.
6:  $\tau_{i}$ = mean(bound_list[-$l$:]).
@@ 4.5 Track the Change of Contribution Bound
7:  $q^{\uparrow}_{i}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D_{i},\tau_{i})$.
38:  (is_up, $\text{count}^{\uparrow}$, $\widetilde{T}^{\uparrow}$)
4= CheckUpdate($q^{\uparrow}_{i}$, $\epsilon^{\uparrow}$, $\text{count}^{\uparrow}$, $k$, $\widetilde{T}^{\uparrow}$, $T^{\uparrow}$).
9:  $q^{\downarrow}_{i}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D_{i},\tau_{i})-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D_{i},\tau_{i}*{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow})$.
510:  (is_down, $\text{count}^{\downarrow}$, $-\widetilde{T}^{\downarrow}$)
6= CheckUpdate($q^{\downarrow}_{i}$, $\epsilon^{\downarrow}$, $\text{count}^{\downarrow}$, $k$, $-\widetilde{T}^{\downarrow}$, $-T^{\downarrow}$).
11:  if is_up == is_down then
12:     $r_{i}=\tau_{i}$.
13:  else if is_up then
14:     $r_{i}=\tau_{i}*{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\uparrow}$.
15:  else if is_down then
16:     $r_{i}=\tau_{i}*{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow}$.
717:  end if
18:  bound_list.append($r_{i}$).
19:  Return $r_{i}$.
Algorithm 4 UpdateBoundSVT
Define ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D_{i},\tau_{i})$ as the number of people who has more than $\tau_{i}$ contributions on day $i$. We can run two parallel SVT to check if there’s a need for updating the contribution bound. One SVT checks if it’s necessary to increase the bound, while the other one checks if it’s better to decrease the bound. Algorithm 3 and Algorithm 4 outline the details of the SVT. The three primary components are discussed below.
Check if the Bound Increases In the first SVT, we use the query function $q^{\uparrow}_{i}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D_{i},\tau_{i})$, here $\tau_{i}$ is the default contribution bound. We want to check if $\tau_{i}*{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\uparrow}$ is a better contribution bound, here ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\uparrow}>1$ is a scaling parameter. If $q^{\uparrow}_{i}$ is larger than a predefined threshold $T^{\uparrow}$, it means that updating $\tau*{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\uparrow}$ as the contribution bound can reduce the bias. Therefore we use $\tau_{i}*s^{\uparrow}$ as the new contribution bound on day $i$. Otherwise, we use $\tau_{i}$ as the contribution bound.
Check if the Bound Decreases In the second SVT, we use the query function $q^{\downarrow}_{i}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D_{i},\tau_{i})-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D_{i},\tau_{i}*{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow})$, here $\tau_{i}$ is the default contribution bound, and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow}<1$. We want to check if $\tau_{i}*{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow}$ is a better contribution bound. If $q_{i}$ is greater than the threshold $-T^{\downarrow}$ (or equivalent, $-q_{i}\leq T^{\downarrow}$), it means that updating $\tau*{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow}$ as the contribution bound can reduce the variance, without introducing too much bias. Therefore we use $\tau_{i}*s^{\downarrow}$ as the new contribution bound on day $i$. Otherwise, we use $\tau_{i}$ as the contribution bound.
Update the Bound For the first $l$ days, we use differentially private quantile algorithm to calculate the private contribution bound as $\tau_{i}$. On $i>l$ we update the default bound as the average of previous $l$ bounds $\tau_{i}=\frac{1}{l}\sum_{j=i-l}^{i-1}\tau_{j}$. If both SVTs attempt to update the bound simultaneously, with one trying to increase it and the other trying to decrease it, we consider this an invalid attempt. In that case, we will use the default bound.
The detailed description of UpdatedBoundSVT is shown in Algorithm 4, we can prove that it satisfies zCDP.
Theorem 4.6.
Algorithm UpdateBoundSVT 4 satisfies $\rho_{3}$-zCDP with $\rho_{3}=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\epsilon$.
$\tau$ 1 2 5 8 10 13 20
Hist[$D,\tau$] 1512 208 53 50 25 12 3
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D,\tau)$ 351 143 90 40 15 3 0
TABLE VII: Statistics of an example dataset $D$. $\tau$ is the contribution bound. Hist$[D,\tau]$ counts the number of people with $\tau$ contributions. ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D,\tau)$ counts the number of people with more than $\tau$ contributions
An Example for SVT To illustrate how SVT can detect changes in user contribution bounds, we use an example dataset $D$. For simplicity and clarity, noise is not added to the queries or thresholds in this demonstration. Table VII shows the statistics of the example dataset $D$. Suppose the default contribution bound is $\tau=10$, the gap parameter is ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\uparrow}=1.5$ and the threshold for SVT is $T^{\uparrow}=100$. We want to check if $\tau*{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\uparrow}=15$ is a better bound. The query value is $q^{\uparrow}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D,\tau)={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D,10)=15$. Because the number of people with contributions above 10 is small (15 is smaller than the threshold $T^{\uparrow}=100$), increasing the default contribution bound can’t reduce enough bias, we don’t want to increase the bound.
Now let’s check if decreasing the default bound is better. The default contribution bound is $\tau=10$, suppose the gap parameter is ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow}=0.5$ and the threshold for SVT is $T^{\downarrow}=100$. We want to check if $\tau*{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow}=5$ is a better bound. The query value is $q^{\downarrow}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D,\tau)-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D,\tau*{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{\downarrow})={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D,10)-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{NumAbove}}}(D,5)=-75$. Because the number of people with contributions between $[5,10]$ is small (75 is smaller than the threshold $T^{\downarrow}=100$), decreasing the default contribution bound can reduce enough variance, we can decrease the bound.
4.6 Extension for Multiple Publishers
@@
Suppose there are $k$ publishers, the data vector for publisher $i$ is ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}_{i}\in\mathbb{R}^{n\times 1}$. Let $\mathbf{X}\in\mathbb{R}^{n\times k}$ be the concatenation of these $k$ data vectors ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}_{i}$. Similar to Section 4.3, the goal is to find a private answer to $\mathbf{Q}\mathbf{X}$. We factorize the workload matrix $\mathbf{Q}=\mathbf{L}\mathbf{B}$, where $\mathbf{B}=diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}})$ and $\mathbf{L}=\mathbf{Q}\mathbf{B}^{-1}$. Let $\mathbf{Z}\sim MVG_{n,k}(0,\mathbf{I}_{n},\mathbf{I}_{k})$, where $MVG_{n,k}(0,\mathbf{I}_{n},\mathbf{I}_{k})$ is a Matrix Gaussian Distribution [12, 53] (which generates $N(0,1)$ noise in each entry of $\mathbf{Z}$), then the mechanism is
$\mathbf{L}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}(\mathbf{X})=\mathbf{L}(\mathbf{B}\mathbf{X}+\mathbf{Z})=\mathbf{Q}\mathbf{X}+\mathbf{L}\mathbf{Z}.$ (10)
@@
Theorem 4.7 (Sensitivity of Mechanism 10).
The sensitivity of the mechanism 10 under the definition of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}$ is
$\Delta_{2}(f)=\max_{(\mathbf{X},\tilde{\mathbf{X}})\in{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{User}}}}}\|\mathbf{B}\mathbf{X}-\mathbf{B}\tilde{\mathbf{X}}\|_{F}=\sqrt{\frac{r_{1}^{2}}{\sigma_{1}^{2}}+\cdots+\frac{r_{n}^{2}}{\sigma_{n}^{2}}}.$ (11)
Then similar to Theorem 4.3, we can show that the mechanism 10 also satisfies zCDP,
@@
Let $\mathbf{Q}\in\mathbb{R}^{m\times n}$ be the input Query Workload matrix. For Mechanism defined in 10, let $\rho_{1}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{budget}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}})$, $\sigma_{1},\sigma_{2},$ $\cdots\sigma_{n}$ be the noise scales, $\mathbf{B}=diag(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}}\cdots,\frac{1}{\sigma_{n}})$, $\mathbf{L}=\mathbf{Q}\mathbf{B}^{-1}$ and $\mathbf{Z}\sim MVG_{n,k}(0,\mathbf{I}_{n},\mathbf{I}_{k})$, then $\mathbf{L}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}(\mathbf{X})$ = $\mathbf{L}(\mathbf{B}\mathbf{X}+\mathbf{Z})=\mathbf{Q}\mathbf{X}+\mathbf{L}\mathbf{Z}$ satisfies $\rho_{1}$-zCDP.
4.7 Combining Together
For Algorithm 2 and Algorithm 4, we convert them from $\epsilon$-DP to $\rho$-zCDP using Theorem 2.4. We prove in the Theorem 4.9 that AdsBPC satisfies $\rho$-zCDP.
Theorem 4.9.
Algorithm AdsBPC 1 satisfies $\rho$-zCDP, with parameter $\rho=\rho_{1}+l*\rho_{2}+\rho_{3}$.
5 Discussion
In AdsBPC, we add independent noise to the data vector with different noise scales. A natural question arises: why not use the matrix mechanism to optimize these noise scales? Our reasons are three-fold. Complexity: Solving the exact matrix mechanism under user-level differential privacy is NP-hard [14]. We compare AdsBPC with two variants of the matrix mechanism that approximate the user-level constraint in our experiments. AdsBPC consistently performs better, while also being simpler—eliminating the need for matrix optimization. Optimality for Small $n$: When the number of days $n$ is small and the objective function is the root mean squared error, we calculated the optimal matrix factorization using the code in [14]. Surprisingly, the resulting strategy matrix is diagonal, yielding the same objective value as AdsBPC. See the Appendix 8.8 for detailed numerical results. Whether AdsBPC achieves optimal noise scales in general remains a question for future research. Flexibility: The matrix mechanism is designed to minimize the sum of variances, limiting it to a single objective function. AdsBPC, on the other hand, offers greater flexibility to tailor to various objectives.
6 Experiments
Throughout the experiments, we denote the length of the data vector as $n=len({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})$. We repeat our experiments 10 times and report the average results. All experiments are conducted on a desktop with an 8-core 16-thread 3.6GHz Intel i9-9900K CPU and 32 GB memory.
6.1 Query and Objective Function
We would like to test the performance of each method under two scenario.
Scenario 1. The advertiser aims to track the trend of conversions over time. Each day, they seek a privacy-preserving estimate of the cumulative conversions up to that day, which corresponds to a prefix-sum query. Additionally, they prioritize accuracy in the final day’s estimate, so the error metric used is the weighted root mean square error (WRMSE), with a higher weight assigned to the last day’s query. In this setting, the initialization of AdsBPC follows Problem 4.3, we set $\gamma_{1}=\cdots=\gamma_{n-1}=1$ and $\gamma_{n}=7$.
Scenario 2. The advertiser is interested in the total number of conversions over the past $K$ days, represented by a sliding window query. Each query is treated as equally important, and the error metric used is the maximum variance. To initialize AdsBPC in this context, we follow Problem 8 and set $v_{1}=\cdots=v_{n}=1$. For comparison with other methods, we scale the noise parameters to ensure that the total privacy budget $\rho$ is consistent across all methods. We then compare the maximum variances across all queries for each method.
6.2 Baseline Methods
We compare our algorithm (named AdsBPC) with other baselines. Unless otherwise specified, we allocate a total privacy budget of $\rho_{total}=1$. Some baseline methods require bounding the global sensitivity. For each dataset, we assume the global sensitivity is given as $GS$. If the global sensitivity is unknown, we can set a reasonably large $GS$ as a threshold. For any user with more than $GS$ records in the dataset, only the first $GS$ records are retained, and any additional records are removed. It is important to note that our algorithm AdsBPC does not require knowledge of $GS$.